# DecoupledESC: Enhancing Emotional Support Generation via Strategy-Response Decoupled Preference Optimization

### v1

@@ Abstract
Recent advances in Emotional Support Conversation (ESC) have improved emotional support generation by fine-tuning Large Language Models (LLMs) via Supervised Fine-Tuning (SFT). However, common psychological errors still persist. While Direct Preference Optimization (DPO) shows promise in reducing such errors through pairwise preference learning, its effectiveness in ESC tasks is limited by two key challenges: (1) Entangled data structure: Existing ESC data inherently entangles psychological strategies and response content, making it difficult to construct high-quality preference pairs; and (2) Optimization ambiguity: Applying vanilla DPO to such entangled pairwise data leads to ambiguous training objectives. To address these issues, we introduce Inferential Preference Mining (IPM) to construct high-quality preference data, forming the IPM-PrefDial dataset. Building upon this data, we propose a Decoupled ESC framework inspired by Gross’s Extended Process Model of Emotion Regulation Gross (2015), which decomposes the ESC task into two sequential subtasks: strategy planning and empathic response generation. Each was trained via SFT and subsequently enhanced by DPO to align with the psychological preference. Extensive experiments demonstrate that our Decoupled ESC framework outperforms joint optimization baselines, reducing preference bias and improving response quality [^1] Our data and code are available at https://anonymous.4open.science/r/DecoupledESC-0E37..
Figure 1: Comparison from Vanilla-SFT to Vanilla-DPO to Decoupled-DPO. Vanilla-SFT lacks negative preference data, leading to high preference bias; Vanilla-DPO uses entangled preference data, causing potential negative optimization (regards PsNr, NsPr as pure negative samples); Decoupled-DPO disentangles strategy and response, effectively reduce bias and psychological errors.
@@ 1 Introduction
Mental health is essential to well-being Prince et al. (2007), yet rising stress and fast-paced life have increased related issues Bor et al. (2014); Brundtland (2000); Paisley and McMahon (2001). According to WHO, 1 in 8 people suffer from mental disorders Organization (2022). Amid a shortage of professionals, this underscores the need for scalable solutions, where Large Language Models (LLMs) offer promising potential.
@@
To enhance the performance of LLMs in Emotional Support Conversation (ESC), prior works Zhang et al. (2024a); Chen et al. (2023) have constructed several large-scale, high-quality dialogue datasets and applied Supervised Fine-Tuning (SFT) to improve model responses. Among them, Liu et al. Liu et al. (2021) built the ESConv dataset based on Hill’s Helping Skills Theory Hill (1999) and filtered out FailedESConv dataset. The ESConv dataset follows a three-phase structure (Exploration → Comfort → Action) and includes eight types of support strategies, each paired with corresponding responses, details are provided in Appendix A and C.1. This structured design significantly enhances a model’s ability to generate empathetic dialogue.
@@
Observation Currently, SFT has become the mainstream approach in the ESC field. However, we observe that models still frequently exhibit common psychological errors Gross (2002); Raskin and Rogers (2005); Stebnicki (2007) during inference, which align with those identified in the FailedESConv dataset (§3.1). In addition, Zhao et al. Zhao et al. (2025) found that SFT’s focus on single gold strategy-response pairs limits adaptability to nuanced contexts, weakening empathetic support. To mitigate this, they use MCTS to collect pairwise preference data linking strategies and responses, and apply Direct Preference Optimization (Vanilla-DPO) to guide the model in choosing appropriate strategies, thereby partially reducing preference bias and improving response quality.
@@
Challenges However, as shown in Figure 1 and  4, our further analysis reveals that the limitations of current work lie not in the SFT or DPO training methods themselves, but rather in two overlooked challenges (§3.2): (1) Entangled data structure: Existing ESC datasets heavily entangle psychological strategies with response content, making it difficult to construct high-quality preference pairs. For instance, penalizing responses with correct strategies but flawed content may degrade data quality. (2) Optimization ambiguity: Applying Vanilla-DPO directly to such entangled data can blur training objectives and even lead to negative optimization outcomes.
@@
Approach To address these issues, we first introduce the Inferential Preference Mining (IPM) method, which automatically constructs preference samples decoupled from strategy-response. Specifically, we use dynamic data routing to route four types of psychological error samples identified from the SFT model’s inference data to the DPO training stage of either strategy planning or response generation, depending on the error type. These samples are then paired with human-annotated ground truth samples to form the Inferential Preference Mining Preference Dialogues (IPM-PrefDial) dataset, containing 21k strategy preference pairs and 11k response preference pairs. This dataset provides disentangled and high-quality supervision signals for two separate DPO models. Building on this, we propose a Decoupled ESC optimization framework, grounded in the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: identification, strategy selection, and implementation, details are provided in Appendix A.1. Accordingly, we explicitly split the ESC task into two subtasks: Strategy Planning (SP) and Response Generation (RG), enabling structured and goal-driven emotional support in dialogue.
@@
  • We analyze common psychological errors in existing SFT paradigms, and introduce the IPM method to construct the IPM-PrefDial, a strategy–response decoupled dataset.
  • We propose a Decoupled ESC framework inspired by Gross’s Extended Process Model of Emotion Regulation, which explicitly splits the ESC task into two subtasks: Strategy Planning and Response Generation, effectively mitigating preference bias and enhancing response quality.
@@
Figure 2: Observation 1: Comparison of common psychological error type proportions among the FailedESConv dataset, Qwen-SFT inference results, and Llama-SFT inference results.
@@ 2 Related Work
Emotional Support Conversation (ESC) aims to alleviate users’ emotional distress through empathetic and supportive responses. Liu et al. Liu et al. (2021) first introduced the concept and built the ESConv dataset with 8 support strategies, 1.3k dialogues. They also released the FailedESConv dataset, containing 196 failed dialogues. Subsequent studies improved ESC systems by enhancing data quality Sun et al. (2021); Qiu et al. (2023); Chen et al. (2023), adding external strategy planners Deng et al. (2024); He et al. (2024, 2025), and incorporating commonsense reasoning Tu et al. (2022); Deng et al. (2023); Ma et al. (2023). Supervised Fine-Tuning (SFT) remains the dominant training paradigm with strong real-world performance (e.g., MeChat Qiu et al. (2023), SweetieChat Ye et al. (2025)). Recently, preference-based methods like Direct Preference Optimization (DPO) Rafailov et al. (2023) have emerged. Zhao et al. Zhao et al. (2025) introduced DPO with MCTS-based data to jointly optimize strategies and responses. However, the fixed coupling limited independent optimization and resulted in lower response quality.
@@
Reinforcement Learning (RL) was initially introduced into LLM training to align with human preferences Ouyang et al. (2022). This approach uses a reward model to guide the optimization of the policy model via the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017). Recently, the Group Relative Policy Optimization (GRPO) algorithm was proposed to enhance model reasoning capabilities Shao et al. (2024), which eliminates the need for a critic model by using within-group rewards as advantages. While these online reinforcement learning methods are effective, they suffer from high computational costs and reliance on accurate reward modeling. As a simpler offline optimization algorithm, DPO optimizes the policy model from pairwise preference data directly without the need for reward modeling. Due to its simplicity and effectiveness, DPO has achieved significant success across multiple domains, including mathematical reasoning, code generation, and recommendation systems Lai et al. (2024); Luo and Yang (2024); Zhang et al. (2024b, 2025); Chen et al. (2024).
@@
3 Preliminary Observations
To investigate the causes of low response quality in the ESC task, we analyzed outputs from six models: Base, SFT, and DPO versions of Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
3.1 Preference Bias and Psychological Errors (Obs 1)
Current base and SFT models (e.g., Qwen-Base, Qwen-SFT) show strong strategy preferences Kang et al. (2024), often overusing fixed strategies and failing to adapt to users’ emotional states. As shown in Figure 3, their strategy distributions diverge significantly from the ground truth.
To further explore the impact of preference bias on response quality, we compared the outputs of Qwen-Base, Qwen-SFT, Llama-Base, and Llama-SFT with the FailedESConv dataset. As shown in Figure 2, common psychological errors Raskin and Rogers (2005); Stebnicki (2007) observed in the SFT-generated responses frequently aligned with those found in FailedESConv, including: (1) Strategy Mismatch, (2) Lack of Empathy, (3) Early Emotion Shift, (4) Template Response, (5) Emotion Misread. The definitions and corresponding examples are detailed in Appendix A.2.
@@ 3.2 Limitations of the DPO Method (Obs 2)
To address Obs 1, a natural approach is to treat filtered failures as negative signals and train with DPO. Prior work Zhao et al. (2025) adopted a vanilla DPO setup that jointly optimizes strategy-response pairs. However, as shown in Figure 1 and Figure 3, Vanilla-DPO relies heavily on the Question strategy and shows a strong preference for it, which fails to significantly reduce preference bias (§6.2).
To investigate the failure of Vanilla-DPO in aligning with human preferences, we conduct a controlled study. We split the preference data based on response quality into two types: (PsPr, PsNr): where the preferred sample has both a positive strategy (Ps) and positive response (Pr), and the non-preferred sample has a positive strategy (Ps) but a negative response (Nr). (PsPr, NsNr): where the non-preferred sample contains both negative strategy (Ns) and negative response (Nr).
We train models using each dataset on Qwen and Llama, and evaluate them on preference bias and strategy preference. As shown in Figure 4, models trained on (PsPr, NsNr) consistently outperform those trained on (PsPr, PsNr). It reduces preference bias and better aligns with diverse strategies. These results show that Vanilla-DPO training with entangled pairs like (PsPr, PsNr) harms strategy learning. This reveals two issues in Vanilla-DPO:
  1. 1.
@@
According to Hill’s Helping Skills Theory Hill (1999) and Gross’s Extended Process Model of Emotion Regulation (EPMER) Gross (2015), strategies should precede response generation and serve as its guidance. In essence, the two are decouplable. However, when strategies and responses are jointly optimized as a single DPO objective, the optimization objective becomes ambiguous. To address this, we propose a decoupled modeling and staged optimization framework for ESC, which separates strategy planning from response generation, enabling more structured and targeted improvements in dialogue quality.
Figure 4: Observation 2: (a) Preference Bias and (b) Strategy Preference across Qwen and Llama models trained on different preference datasets.
4 Datasets
4.1 Preference Dataset Construction
Inferential Preference Mining.
Standard SFT approaches focus on expert demonstrations and neglect failure cases, limiting the model’s ability to learn from mistakes. This can result in preference bias and the generation of low-quality responses.
To address these issues, we propose an Inferential Preference Mining (IPM) method to collect high-quality preference data, along with a dynamic data routing mechanism that identifies psy-errors and routes data to targeted DPO training phases.
Specifically, we first perform inference using Qwen-SFT and Llama-SFT models to generate inference data, which are paired with the corresponding ground truth to form candidate preference pairs. Subsequently, we apply a dynamic data routing mechanism to filter these pairs based on the four types of common psy-errors observed in Obs1 (§3.1). This process yields the final IPM-PrefDial dataset. Concretely, for each subtask:
For the Strategy Planner (SP), we identify suboptimal strategies $s_{r}$ exhibiting the Strategy Mismatch psy-error, pair them with the gold strategies $s_{c}$ and context $c$, and construct the $D_{\text{SP-dpo}}$ dataset, which is routed to the SP’s DPO training phase:
For the Response Generator (RG), the SFT model generates multiple responses based on the gold strategy. Suboptimal responses $a_{r}$ exhibiting typical psychological errors, namely Lack of Empathy, Early Emotion Shift, or Template Response, are identified via the dynamic data routing. These are then paired with the corresponding gold responses $a_{c}$ to construct the $D_{\text{RG-dpo}}$ dataset, which is routed to the RG’s DPO training phase:
Figure 5: Comparison between previous vanilla SFT training paradigm and our proposed Decoupled ESC framework.
@@ 4.1 Preference Dataset Construction
Table 1: Statistics of the ESConv and IPM-PrefDial Datasets.

### v2

@@ Abstract
Recent advances in Emotional Support Conversation (ESC) have improved emotional support generation by fine-tuning Large Language Models (LLMs) via Supervised Fine-Tuning (SFT). However, common psychological errors still persist. While Direct Preference Optimization (DPO) shows promise in reducing such errors through pairwise preference learning, its effectiveness in ESC tasks is limited by two key challenges: (1) Entangled data structure: Existing ESC data inherently entangles psychological strategies and response content, making it difficult to construct high-quality preference pairs; and (2) Optimization ambiguity: Applying vanilla DPO to such entangled pairwise data leads to ambiguous training objectives. To address these issues, we introduce Inferential Preference Mining (IPM) to construct high-quality preference data, forming the IPM-PrefDial dataset. Building upon this data, we propose a Decoupled ESC framework inspired by Gross’s Extended Process Model of Emotion Regulation, which decomposes the ESC task into two sequential subtasks: strategy planning and empathic response generation. Each was trained via SFT and subsequently enhanced by DPO to align with the psychological preference. Extensive experiments demonstrate that our Decoupled ESC framework outperforms baselines, reducing preference bias and improving response quality [^1] Our data and code are available at https://github.com/Zc0812/DecoupledESC..
Figure 1: Comparison from Vanilla-SFT to Vanilla-DPO to Decoupled-DPO. Vanilla-SFT lacks negative preference data, leading to high preference bias; Vanilla-DPO uses coupled preference data, causing potential negative optimization (regards PsNr, NsPr as pure negative samples); Decoupled-DPO decouples strategy and response, effectively reducing bias and psychological errors.
@@ 1 Introduction
Mental health is essential to well-being Prince et al. (2007), yet rising stress and fast-paced life have increased related issues Bor et al. (2014); Brundtland (2000); Paisley and McMahon (2001); Ma et al. (2023). According to WHO, 1/8 people suffer from mental disorders Organization (2022). Amid a shortage of professionals, this underscores the need for scalable solutions, where Large Language Models (LLMs) offer promising potential.
@@
To enhance the performance of LLMs in Emotional Support Conversation (ESC), prior works Zhang et al. (2024); Chen et al. (2023) have constructed several large-scale, high-quality dialogue datasets and applied Supervised Fine-Tuning (SFT) to improve model responses. Among them, Liu et al. Liu et al. (2021) built the ESConv dataset based on Hill’s Helping Skills Theory Hill (1999) and filtered out FailedESConv dataset. The ESConv dataset follows a three-phase structure (Exploration → Comfort → Action) and includes eight types of support strategies, each paired with corresponding responses, details are provided in Appendix A and C.1. This structured design significantly enhances a model’s ability to generate empathetic dialogue.
@@
Observation Currently, SFT has become the mainstream approach in the ESC field. However, we observe that models still frequently exhibit common psychological errors Raskin and Rogers (2005); Stebnicki (2007) during inference, which align with those identified in the FailedESConv dataset (Obs 1). In addition, Zhao et al. Zhao et al. (2025) found that SFT’s focus on single gold strategy-response pairs limits adaptability to nuanced contexts, weakening empathetic support. To mitigate this, they use Monte Carlo Tree Search (MCTS) to collect pairwise preference data linking strategies and responses, and apply Direct Preference Optimization (Vanilla-DPO) to guide the model in choosing appropriate strategies, thereby partially reducing preference bias and improving response quality.
@@
Challenges However, as shown in Figure 1 and  4, our further analysis reveals that the limitations of current work lie not in the SFT or DPO training methods themselves, but rather in two overlooked challenges (Obs 2): (1) Entangled data structure: Existing ESC datasets heavily entangle psychological strategies with response content, making it difficult to construct high-quality preference pairs. For instance, penalizing responses with correct strategies but flawed content may degrade data quality. (2) Optimization ambiguity: Applying Vanilla-DPO directly to such entangled data can blur training objectives and even lead to negative optimization outcomes.
@@
Approach To address these issues, we first introduce the Inferential Preference Mining (IPM) method, which automatically constructs preference samples decoupled from strategy-response. Specifically, we use dynamic data routing to route four types of psychological error samples identified from the SFT model’s inference data to the DPO training stage of either strategy planning or response generation, depending on the error type. These samples are then paired with human-annotated ground truth samples to form the Inferential Preference Mining Preference Dialogues (IPM-PrefDial) dataset, containing 2121k strategy preference pairs and 1111k response preference pairs. This dataset provides disentangled and high-quality supervision signals for two separate DPO models. Building on this, we propose a Decoupled ESC optimization framework (DecoupledESC), grounded in the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: (1) Identification → (2) Strategy Selection → (3) Implementation, details are provided in Appendix A.1. Accordingly, we explicitly split the ESC task into two subtasks: Strategy Planning (SP) and Response Generation (RG).
@@
  • We analyze common psychological errors in existing SFT paradigms and introduce Inferential Preference Mining (IPM) and Expert-Guided ICL Annotation to construct IPM-PrefDial, a decoupled strategy–response dataset.
  • We propose a Decoupled ESC framework, which explicitly splits the ESC task into two subtasks: Strategy Planning and Response Generation, effectively mitigating preference bias and enhancing response quality.
@@
Figure 2: Comparison of common psychological error type proportions among the FailedESConv dataset, Qwen-SFT inference results, and Llama-SFT inference results. Other Error refers to non-psychological errors.
Subset Qu RP RF Sd AR PS In Ot
GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$
1 110 135 1.41 40 25 0.57 43 38 0.82 39 50 1.43 85 50 0.55 77 91 1.29 26 19 0.69 80 92 1.25
2 104 122 1.32 35 18 0.51 41 43 1.11 48 42 0.90 74 56 0.78 73 95 1.46 28 20 0.76 97 104 1.15
3 97 120 1.42 29 30 1.04 47 37 0.79 37 40 1.13 79 54 0.66 83 95 1.24 35 20 0.57 93 104 1.16
4 93 135 1.78 27 13 0.45 40 32 0.81 40 42 1.19 89 49 0.51 77 110 1.68 32 18 0.57 102 101 1.02
Table 1: Analysis of Strategy Distribution and Preference Bias ($\mathcal{B}$) Across Four Randomly Sampled Subsets. This expanded layout clarifies the relationship between Ground Truth (GT), Predictions (Pred), and the resulting Bias for each strategy (defined in Appendix A.4).
@@ 2 Related Work
Emotional Support Conversation (ESC) aims to alleviate users’ emotional distress through empathetic and supportive responses. Liu et al. Liu et al. (2021) first introduced the concept and built the ESConv dataset with 8 support strategies, 1.3k dialogues. They also released the FailedESConv dataset, containing 196 failed dialogues. Subsequent studies improved ESC systems by enhancing data quality Sun et al. (2021); Qiu et al. (2024); Chen et al. (2023), adding external strategy planners Deng et al. (2024); He et al. (2024, 2025), and incorporating commonsense reasoning Tu et al. (2022); Deng et al. (2023); Luo and Yang (2024). Supervised Fine-Tuning (SFT) remains the dominant training paradigm with strong real-world performance (e.g., MeChat Qiu et al. (2024), SweetieChat Ye et al. (2025)). Recently, preference-based methods like Direct Preference Optimization (DPO) Rafailov et al. (2023) have emerged. Zhao et al. Zhao et al. (2025) introduced DPO with MCTS-based data to jointly optimize strategies and responses. However, the fixed coupling limited independent optimization and resulted in lower response quality.
@@
Reinforcement Learning (RL) was initially introduced into LLM training to align with human preferences Ouyang et al. (2022). This approach uses a reward model to guide the optimization of the policy model via the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017). Recently, the Group Relative Policy Optimization (GRPO) algorithm was proposed to enhance model reasoning capabilities Shao et al. (2024), which eliminates the need for a critic model by using within-group rewards as advantages. While these online reinforcement learning methods are effective, they suffer from high computational costs and reliance on accurate reward modeling. As a simpler offline optimization algorithm, DPO optimizes the policy model from pairwise preference data directly without the need for reward modeling. Due to its simplicity and effectiveness, DPO has achieved significant success across multiple domains, including mathematical reasoning, code generation, and recommendation systems Lai et al. (2024); Zhang et al. (2025a, b); Chen et al. (2024).
3 Preliminary Observations
To investigate the causes of low response quality in the ESC task, we analyzed outputs from six models: Base, SFT [^2] Trained on ESConv Liu et al. (2021) datasets.
github.com/thu-coai/Emotional-Support-Conversation
, and DPO versions of Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
3.1 Preference Bias and Psy-Errors (Obs 1)
Current Base and SFT models (e.g., Qwen-Base, Qwen-SFT) show strong strategy preferences Kang et al. (2024), often overusing fixed strategies and failing to adapt to users’ emotional states. As shown in Figure 3, the Base and SFT models show different levels of divergence from the ground truth in strategy distributions.
Additionally, to verify that the observed strategy preference is an intrinsic model bias, not an artifact of sampling variance, we performed a validation on four random subsets of the test set. As shown in Table 1, the results reveal a consistent strategy bias across all subsets. Despite fluctuations in ground truth counts, the preference bias remained stable, strongly suggesting it is an intrinsic characteristic of the model’s decision-making process.
To further explore the impact of preference bias on response quality, we compared the outputs of Qwen-Base, Qwen-SFT, Llama-Base, and Llama-SFT with the FailedESConv dataset. As shown in Figure 2, common psychological errors Raskin and Rogers (2005); Stebnicki (2007) observed in the SFT-generated responses frequently aligned with those found in FailedESConv [^3] We have manually filtered out samples classified as FailedESConv due to non-psychological errors, such as incomplete or repetitive dialogues., including: (1) Strategy Mismatch, (2) Lack of Empathy, (3) Early Emotion Shift, (4) Template Response, (5) Emotion Misread. The definitions and corresponding examples are detailed in Appendix A.2.
@@ 3.1 Preference Bias and Psy-Errors (Obs 1)
Figure 4: (a) Preference Bias and (b) Strategy Preference across Qwen and Llama models trained on different preference datasets.
@@ 3.2 Limitations of the DPO Method (Obs 2)
To address Obs 1, a natural approach is to treat filtered failures as negative signals and train with DPO. Prior work Zhao et al. (2025) adopted a vanilla DPO setup that jointly optimizes strategy-response pairs. However, as shown in Figure 1 and Figure 3, Vanilla-DPO relies heavily on the Question strategy and shows a strong preference for it, which fails to significantly reduce preference bias (see section Results).
To investigate the failure of Vanilla-DPO in aligning with human preferences, we conduct a controlled study. We split the preference data into two types: ① (PsPr, PsNr): where the preferred sample has both a positive strategy (Ps) and positive response (Pr), and the non-preferred sample has a positive strategy (Ps) but a negative response (Nr). ② (PsPr, NsNr): where the non-preferred sample contains both negative strategy (Ns) and negative response (Nr).
We train models using each dataset on Qwen and Llama, and evaluate them on preference bias and strategy preference. As shown in Figure 4 (a) and (b), models trained on consistently outperform those trained on . It reduces preference bias and better aligns with diverse strategies.
These results show that Vanilla-DPO training with entangled pairs like harms strategy learning. This reveals two issues in Vanilla-DPO:
  1. (1)
@@
According to Hill’s Helping Skills Theory Hill (1999) and Gross’s Extended Process Model of Emotion Regulation (EPMER) Gross (2015), strategies should precede response generation and serve as its guidance. In essence, the two are decouplable. Inspired by this, we propose a decoupled modeling and staged optimization framework for ESC, which separates strategy planning from response generation, enabling more structured and targeted improvements in dialogue quality.
@@
Table 2: Statistics of the ESConv and IPM-PrefDial Datasets. Char: Character, Pref-Pairs: Preference Pairs.
4 Datasets
4.1 Preference Dataset Construction
Expert-Guided ICL Annotation.
Due to the large scale of the dataset, relying solely on human experts for psychological error annotation is highly time- and labor-intensive. Therefore, we adopt an In-Context Learning (ICL) Brown et al. (2020) approach to guide LLM-based classification. Specifically, we engaged 3 professional psychologists to annotate representative examples of 5 common psychological errors, along with detailed explanations. These expert-labeled instances were then used as ICL prompts, enabling the LLM to perform classification in alignment with expert standards.
Inferential Preference Mining.
The absence of failure-aware learning in standard SFT contributes to psychological errors observed in Obs 1. To address these issues, we propose Inferential Preference Mining (IPM) for collecting high-quality preference data, and an Expert-Guided ICL approach that leverages LLMs’ in-context learning to identify psychological errors.
Specifically, we first use the Qwen-SFT and Llama-SFT models to generate responses on the ESConv dataset. These responses are then paired with the corresponding gold responses to construct candidate preference pairs. To ensure consistency between the LLM and psychological expert annotators, we apply an Expert-Guided ICL Annotation process that prompts the LLM to filter out pairs exhibiting four common types of psychological errors. The remaining high-quality pairs constitute the IPM-PrefDial dataset.
More concretely, for the Strategy Planner (SP), we pair suboptimal strategies $s_{r}$ exhibiting the psychological error (1) Strategy Mismatch with gold strategies $s_{c}$ and context $c$ to form $D_{\text{SP-dpo}}$:
For the Response Generator (RG), the SFT model produces multiple responses based on the gold strategy. Suboptimal responses $a_{r}$ exhibiting the psychological errors (2) Lack of Empathy, (3) Early Emotion Shift, or (4) Template Response are identified via Expert-Guided ICL and paired with gold responses $a_{c}$ to form $D_{\text{RG-dpo}}$:
Figure 5: Comparison between previous vanilla SFT training paradigm and our proposed Decoupled ESC framework. The Decoupled ESC first undergoes SFT initialization, followed by DPO training using the IPM-PrefDial dataset.
[…]
Qwen 2.5-7B- Instruct Chosen 3.82 3.52 3.20 2.95
Rejected 3.65 3.09 2.41 2.32
Improve ($\uparrow$) 4.66% 13.92% 32.78% 27.16%
Llama 3.1-8B- Instruct Chosen 3.99 3.74 3.33 3.09
Rejected 3.93 3.21 2.40 2.39
Improve ($\uparrow$) 1.53% 16.52% 38.75% 29.29%